. If Bubba’s sells 52,000 hats, what is its margin of safety (MOS) in dollars and MOS ratio? 4. Bubba is considering the elimination of sales commissions completely and increasing salaries by $357,000 annually. What would be the new breakeven point in units? What would be the before-tax income or loss if 39,000 hats are sold with the new salary plan? rev: 02_10_2020_QC_CS-200042

Bubba’s Western Wear is a western hat retailer in Lubbock, Texas. Although Bubba’s carries numerous styles of western hats, each hat has approximately the same price and invoice (purchase) cost, as shown in the following table. Sales personnel receive a commission to encourage them to be more aggressive in their sales efforts. Currently, the Lubbock economy is really humming, and sales growth at Bubba’s has been great. The business is very competitive, however, and Bubba, the owner, has relied on his knowledgeable and courteous staff to attract and retain customers who otherwise might go to other western wear stores. Because of the rapid growth in sales, Bubba is also finding the management of certain aspects of the business more difficult, such as restocking inventory and hiring and training new salespeople.

Sales price $ 80.00
Per unit variable expenses
Purchase cost 23.50
Sales commissions 31.50
Total per unit variable costs $ 55.00
Total annual fixed expenses
Advertising $ 297,500
Rent 248,000
Salaries 397,000
Total fixed expenses $ 942,500

Required:

1. Calculate the annual breakeven point, both in terms of units and in terms of sales dollars.

2. If Bubba’s sells 39,000 hats, what is its before-tax income or loss? Support your answer by preparing a contribution income statement.

3. If Bubba’s sells 52,000 hats, what is its margin of safety (MOS) in dollars and MOS ratio?

4. Bubba is considering the elimination of sales commissions completely and increasing salaries by $357,000 annually. What would be the new breakeven point in units? What would be the before-tax income or loss if 39,000 hats are sold with the new salary plan?
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

Answer Part 1) Annual Break even point = Fixed cost / contribution margin per unit =942500/(80-55) = 37700 units break even d

Part 3) Margin of safety in dollars = (total sales in units - break even sales in units)*selling price =(52000-37700)*80 = $